# A Position Paper on the Automatic Generation of Machine Learning Leaderboards

### v1

@@ Appendix F An Overview of Experimental Results
We have compiled all the results we could find in the literature where researchers experiment with extracting leaderboard tuples and entities, evaluating these extractions using micro, partial micro, or macro precision, recall, and F1 scores. Table LABEL:tab:experiments2 presents an overview of these experiments. This table highlights the complexity of comparing different results due to the diversity of problem framing (e.g. closed versus open domain), datasets and metrics.. We omitted details on how the scores were averaged (e.g., across papers or leaderboards), as this information is often not reported in many studies. These differences in averaging methods also complicate direct comparisons between works. Please note that there may be additional subtle variations in the experimental setup that are not captured in the table, which could prevent a fair comparison.
| Micro | Macro | Part. Micro | ||||||||||
| Reported In | P | R | F1 | P | R | F1 | P | F1 | Dataset | Method | Experimental Setup | |
| Results of Extracting Task, Dataset, Metric for Closed Domain Problem Framing | ||||||||||||
| Hou2019Leaderboards | 60.2 | 73.1 | 66.0 | 54.1 | 65.9 | 56.6 | NLP-TDMS-v1 | TDMS-IE | ||||
| Hou2019Leaderboards | 29.4 | 42.0 | 34.6 | 24.9 | 43.6 | 28.1 | NLP-TDMS-v1 | EL† | ||||
| Hou2019Leaderboards | 56.8 | 23.8 | 33.6 | 56.8 | 30.9 | 37.3 | NLP-TDMS-v1 | MLC† | ||||
| Hou2019Leaderboards | 16.8 | 7.8 | 10.6 | 8.1 | 6.4 | 6.9 | NLP-TDMS-v1 | SM† | ||||
| Hou2019Leaderboards | 60.8 | 76.8 | 67.8 | 62.5 | 75.2 | 65.3 | NLP-TDMS-v2 | TDMS-IE | ||||
| Hou2019Leaderboards | 24.3 | 36.3 | 29.1 | 18.1 | 31.8 | 20.5 | NLP-TDMS-v2 | EL† | ||||
| Hou2019Leaderboards | 42.0 | 20.9 | 27.9 | 42.0 | 23.1 | 27.8 | NLP-TDMS-v2 | MLC† | ||||
| Hou2019Leaderboards | 36.0 | 19.6 | 25.4 | 31.8 | 30.6 | 31.0 | NLP-TDMS-v2 | SM† | ||||
| Hou2019Leaderboards | 68.6 | 40.3 | 50.8 | 29.6 | 29.1 | 28.1 | NLP-TDMS-v2 | TDMS-IE | TAE# | |||
| Hou2019Leaderboards | 50.0 | 23.7 | 32.2 | 20.8 | 20.1 | 19.4 | NLP-TDMS-v2 | TDMS-IE | TAT# | |||
| Hou2019Leaderboards | 47.9 | 14.2 | 21.9 | 11.3 | 11.3 | 10.7 | NLP-TDMS-v2 | TDMS-IE | TA# | |||
| Kardas2020AXCELL | 65.8 | 58.5 | 61.9 | 56.0 | 55.8 | 54.1 | NLP-TDMS-v3 | AxCell | ||||
| Kardas2020AXCELL | 53.4 | 66.3 | 59.2 | 57.1 | 66.1 | 58.5 | NLP-TDMS-v3 | TDMS-IE | ||||
| Kardas2020AXCELL | 67.8 | 47.8 | 56.1 | 47.9 | 46.4 | 43.5 | PwC-LB-v1 | AxCell | ||||
| Kabongo2021Mining | 76.4 | 66.4 | 71.1 | 63.5 | 64.1 | 61.4 | NLP-TDMS-v1 | ORKG-TDM | XLNet | |||
| Kabongo2021Mining | 65.3 | 73.1 | 69.0 | 57.6 | 68.7 | 60.1 | NLP-TDMS-v1 | ORKG-TDM | SciBERT | |||
| Kabongo2021Mining | 79.5 | 57.6 | 66.8 | 59.0 | 55.4 | 54.7 | NLP-TDMS-v1 | ORKG-TDM | BERT | |||
| Kabongo2021Mining | 77.1 | 70.9 | 73.9 | 71.7 | 73.9 | 70.6 | NLP-TDMS-v2 | ORKG-TDM | XLNet | |||
| Kabongo2021Mining | 79.6 | 63.3 | 70.5 | 68.1 | 67.5 | 65.5 | NLP-TDMS-v2 | ORKG-TDM | BERT | |||
| Kabongo2021Mining | 65.7 | 76.8 | 70.8 | 65.7 | 77.2 | 68.3 | NLP-TDMS-v2 | ORKG-TDM | SciBERT | |||
| Kabongo2021Mining | 95.1 | 92 | 93.5 | 92.3 | 93.5 | 91.7 | ORKG-PwC-v1 | ORKG-TDM | XLNet TAET# | |||
| Kabongo2021Mining | 93.5 | 93.2 | 93.3 | 90.5 | 94.4 | 91.2 | ORKG-PwC-v1 | ORKG-TDM | XLNet TAT# | |||
| Kabongo2021Mining | 95.0 | 90.5 | 92.7 | 91.6 | 93.1 | 91.2 | ORKG-PwC-v1 | ORKG-TDM | XLNet# | |||
| Kabongo2021Mining | 95.7 | 88.3 | 91.8 | 91.7 | 92.1 | 90.8 | ORKG-PwC-v1 | ORKG-TDM | BERT | |||
| Kabongo2021Mining | 94.2 | 89 | 91.5 | 89.2 | 91.5 | 89.2 | ORKG-PwC-v1 | ORKG-TDM | XLNet TAE# | |||
| Kabongo2021Mining | 94.4 | 87.6 | 90.9 | 89.7 | 91.4 | 89.4 | ORKG-PwC-v1 | ORKG-TDM | SciBERT | |||
| Kabongo2021Mining | 92.6 | 90 | 91.3 | 88.6 | 92.9 | 89.4 | ORKG-PwC-v1 | ORKG-TDM | XLNet TA# | |||
| Kabongo2021Mining | 94.9 | 91.2 | 93.0 | 92.8 | 94.8 | 92.8 | ORKG-PwC-v2 | ORKG-TDM | XLNet | |||
| Kabongo2021Mining | 95.5 | 89.1 | 92.1 | 92.8 | 93.9 | 92.4 | ORKG-PwC-v2 | ORKG-TDM | BERT | |||
| Kabongo2021Mining | 94.1 | 88.5 | 91.2 | 90.9 | 93.4 | 91.1 | ORKG-PwC-v2 | ORKG-TDM | SciBERT | |||
| Kabongo2023ORKG | 95.2 | 92.2 | 93.6 | 91.5 | 93.3 | 91.3 | ORKG-PwC-v5 | ORKG-LB | BigBERT | |||
| Kabongo2023ORKG | 94.8 | 93.9 | 94.3 | 91.3 | 94.4 | 91.8 | ORKG-PwC-v5 | ORKG-LB | BERT | |||
| Kabongo2023ORKG | 94.8 | 93.9 | 94.3 | 91.3 | 94.4 | 91.8 | ORKG-PwC-v5 | ORKG-LB | SciBERT | |||
| Kabongo2023ORKG | 95.4 | 93.9 | 94.7 | 93.2 | 95.7 | 93.5 | ORKG-PwC-v6 | ORKG-LB | BERT | |||
| Kabongo2023ORKG | 95.4 | 91.1 | 93.2 | 92.6 | 94.3 | 92.2 | ORKG-PwC-v6 | ORKG-LB | SciBERT | |||
| Kabongo2023ORKG | 93.2 | 94.9 | 93.0 | 95.7 | 92.4 | 94.0 | ORKG-PwC-v6 | ORKG-LB | BigBERT | |||
| Kabongo2023ORKG | 95.1 | 94.6 | 94.8 | 93.1 | 96.4 | 93.7 | ORKG-PwC-v6 | ORKG-LB | XLNet | |||
| Kabongo2023ORKG | 95.4 | 88.0 | 91.5 | 91.2 | 92.3 | 90.6 | ORKG-PwC-v3 | ORKG-LB | BERT | |||
| Kabongo2023ORKG | 93.7 | 86.0 | 89.7 | 89.4 | 91.7 | 89.2 | ORKG-PwC-v3 | ORKG-LB | SciBERT | |||
| Kabongo2023ORKG | 93.6 | 85.3 | 89.3 | 87.5 | 88.7 | 86.6 | ORKG-PwC-v3 | ORKG-LB | BigBird | |||
| Kabongo2023ORKG | 94.9 | 91.2 | 93.0 | 91.9 | 94.4 | 92.0 | ORKG-PwC-v4 | ORKG-LB | XLNet | |||
| Kabongo2023ORKG | 96.0 | 90.0 | 92.9 | 93.5 | 94.2 | 92.8 | ORKG-PwC-v4 | ORKG-LB | BERT | |||
| Kabongo2023ORKG | 94.6 | 88.6 | 91.5 | 91.7 | 93.9 | 91.6 | ORKG-PwC-v4 | ORKG-LB | SciBERT | |||
| Kabongo2023ORKG | 94.6 | 87.2 | 90.7 | 90.7 | 91.6 | 89.7 | ORKG-PwC-v4 | ORKG-LB | BigBird | |||
| Kabongo2023ZeroShot | 9.2 | 78.1 | 16.5 | 14.3 | 86.6 | 21.9 | ORKG-PwC-v7* | ORKG-TDM | XLNet | |||
| Kabongo2023ZeroShot | 14.1 | 72.9 | 23.6 | 20.1 | 83.4 | 28.9 | ORKG-PwC-v7* | ORKG-TDM | BERT | |||
| Kabongo2023ZeroShot | 10.4 | 81.7 | 18.4 | 16.2 | 89 | 24.4 | ORKG-PwC-v7* | ORKG-TDM | BERT | |||
| Kabongo2023ZeroShot | 10.1 | 76.8 | 17.8 | 14.9 | 86.4 | 22.7 | ORKG-PwC-v7* | ORKG-TDM | XLNet | |||
| sahinuc2024efficient | 55.1 | 25.8 | 35.1 | SciLead | AxCell | |||||||
| sahinuc2024efficient | 40.7 | 39.5 | 40.1 | SciLead | TDMR-PR | Llama 2+CS | ||||||
| sahinuc2024efficient | 35.9 | 34.9 | 35.4 | SciLead | TDMR-PR | Llama 2 | ||||||
| sahinuc2024efficient | 58.4 | 52.1 | 55.1 | SciLead | TDMR-PR | Mixtral+CS | ||||||
| sahinuc2024efficient | 55.7 | 48.8 | 51.0 | SciLead | TDMR-PR | Mixtral | ||||||
| sahinuc2024efficient | 62.0 | 58.1 | 60.0 | SciLead | TDMR-PR | Llama 3+CS | ||||||
| sahinuc2024efficient | 77.1 | 72.6 | 74.8 | SciLead | TDMR-PR | Llama 3 | ||||||
| sahinuc2024efficient | 69.0 | 63.8 | 66.3 | SciLead | TDMR-PR | GPT-4+CS | ||||||
| sahinuc2024efficient | 75.3 | 70.4 | 72.8 | SciLead | TDMR-PR | GPT-4 | ||||||
| Results of Extracting Task, Dataset, Metric for Open Domain Problem Framing | ||||||||||||
| Yang2022TELIN | 68.2 | 45.3 | 56.5 | 49.7 | 43.1 | 42.5 | PwC-LB-v2 | TELIN | ||||
| Results of Extracting Task, Dataset, Metric for Hybrid Domain Problem Framing | ||||||||||||
| sahinuc2024efficient | 27.23 | 22.99 | 24.93 | SciLead | TDMR-PR | Llama 2 | ||||||
| sahinuc2024efficient | 27.89 | 24.48 | 26.07 | SciLead | TDMR-PR | Mixtral | ||||||
| sahinuc2024efficient | 50.75 | 45.30 | 47.87 | SciLead | TDMR-PR | Llama 3 | ||||||
| sahinuc2024efficient | 56.08 | 51.89 | 53.90 | SciLead | TDMR-PR | GPT-4 | ||||||
| Results of Extracting Task, Dataset, Metric, Score for Closed Domain Problem Framing | ||||||||||||
| Hou2019Leaderboards | 10.8 | 13.1 | 11.8 | 9.3 | 11.8 | 9.9 | NLP-TDMS-v1 | TDMS-IE | ||||
| Hou2019Leaderboards | 3.8 | 1.8 | 2.4 | 1.3 | 1.0 | 1.1 | NLP-TDMS-v1 | SM† | ||||
| Hou2019Leaderboards | 6.8 | 2.9 | 4.0 | 6.8 | 6.1 | 6.2 | NLP-TDMS-v1 | MLC† | ||||
| Kardas2020AXCELL | 27.4 | 24.4 | 25.8 | 20.2 | 20.6 | 19.7 | NLP-TDMS-v3 | AxCell | ||||
| Kardas2020AXCELL | 6.8 | 8.4 | 7.5 | 8.6 | 9.5 | 8.8 | NLP-TDMS-v3 | TDMS-IE | ||||
| Kardas2020AXCELL | 37.4 | 23.2 | 28.7 | 24.0 | 21.8 | 21.1 | PwC-LB-v1 | AxCell | ||||
| sahinuc2024efficient | 32.59 | 13.67 | 19.26 | SciLead | AxCell | |||||||
| sahinuc2024efficient | 10.06 | 21.59 | 13.73 | SciLead | TDMR-PR | Llama 2+CS | ||||||
| sahinuc2024efficient | 9.63 | 15.25 | 11.81 | SciLead | TDMR-PR | Llama 2 | ||||||
| sahinuc2024efficient | 26.54 | 24.61 | 25.54 | SciLead | TDMR-PR | Mixtral+CS | ||||||
| sahinuc2024efficient | 24.66 | 21.73 | 23.10 | SciLead | TDMR-PR | Mixtral | ||||||
| sahinuc2024efficient | 23.22 | 29.54 | 26.00 | SciLead | TDMR-PR | Llama 3+CS | ||||||
| sahinuc2024efficient | 27.11 | 35.60 | 30.78 | SciLead | TDMR-PR | Llama 3 | ||||||
| sahinuc2024efficient | 49.82 | 48.71 | 49.26 | SciLead | TDMR-PR | GPT-4+CS | ||||||
| sahinuc2024efficient | 56.02 | 54.53 | 55.27 | SciLead | TDMR-PR | GPT-4 | ||||||
| Results of Extracting Task, Dataset, Metric, Score for Open Domain Problem Framing | ||||||||||||
| Yang2022TELIN | 38.3 | 20.8 | 26.3 | 26.6 | 19.2 | 21.3 | PwC-LB-v2 | TELIN | ||||
| Results of Extracting Task, Dataset, Metric, Score for Hybrid Domain Problem Framing | ||||||||||||
| sahinuc2024efficient | 4.17 | 9.89 | 5.87 | SciLead | TDMR-PR | Llama 2 | ||||||
| sahinuc2024efficient | 14.65 | 12.27 | 13.35 | SciLead | TDMR-PR | Mixtral | ||||||
| sahinuc2024efficient | 15.70 | 18.75 | 17.09 | SciLead | TDMR-PR | Llama 3 | ||||||
| sahinuc2024efficient | 40.60 | 39.56 | 40.07 | SciLead | TDMR-PR | GPT-4 | ||||||
| sahinuc2024efficient | 51.01 | 51.03 | 51.02 | SciLead | TDMR-PR | GPT-4 FS | ||||||
| Results of Extracting Task, Dataset, Metric, Method for Closed Domain Problem Framing | ||||||||||||
| Jain2020SCIREX | 0.48 | 0.89 | 0.62 | SciREX | TDMS-IE | |||||||
| Results of Extracting Task, Dataset, Metric, Method for Open Domain Problem Framing | ||||||||||||
| Jain2020SCIREX | 0.53 | 0.72 | 0.61 | SciREX | SciREX-IE | |||||||
| Results of Extracting Task for Closed Domain Problem Framing | ||||||||||||
| Kardas2020AXCELL | 70.6 | 57.3 | 63.3 | 60.7 | 62.6 | 59.7 | PwC-LB-v1 | AxCell | ||||
| Kabongo2021Mining | 97.4 | 93.6 | 95.5 | 93.7 | 94.8 | 93.6 | ORKG-PwC-v1 | ORKG-TDM | XLNet | |||
| Kabongo2023ORKG | 96.8 | 95.9 | 96.4 | 94.3 | 97.2 | 95.0 | ORKG-PwC-v6 | ORKG-LB | XLNet | |||
| Kabongo2023ORKG | 96.8 | 95.9 | 96.4 | 94.3 | 97.2 | 95.0 | ORKG-PwC-v4 | ORKG-LB | XLNet | |||
| sahinuc2024efficient | 68.98 | 58.52 | 63.32 | SciLead | AxCell | |||||||
| sahinuc2024efficient | 59.83 | 67.20 | 63.30 | SciLead | TDMR-PR | Llama 2+CS | ||||||
| sahinuc2024efficient | 55.45 | 60.74 | 57.97 | SciLead | TDMR-PR | Llama 2 | ||||||
| sahinuc2024efficient | 86.27 | 91.99 | 89.04 | SciLead | TDMR-PR | Mixtral+CS | ||||||
| sahinuc2024efficient | 86.85 | 89.74 | 88.27 | SciLead | TDMR-PR | Mixtral | ||||||
| sahinuc2024efficient | 85.69 | 90.85 | 88.19 | SciLead | TDMR-PR | Llama 3+CS | ||||||
| sahinuc2024efficient | 87.33 | 92.17 | 89.68 | SciLead | TDMR-PR | Llama 3 | ||||||
| sahinuc2024efficient | 90.70 | 90.77 | 90.73 | SciLead | TDMR-PR | GPT-4+CS | ||||||
| sahinuc2024efficient | 90.62 | 91.10 | 90.86 | SciLead | TDMR-PR | GPT-4 | ||||||
[…]

### v2

@@ Appendix G An Overview of Experimental Results
We have compiled all the results we could find in the literature where researchers experiment with extracting leaderboard tuples and entities, evaluating these extractions using micro, partial micro, or macro precision, recall, and F1 scores. Tables 8 - 15 present an overview of these experiments. These tables highlight the complexity of comparing different results due to the diversity of problem framing (e.g. closed versus open domain), datasets and metrics. We omitted details on how the scores were averaged (e.g., across papers or leaderboards), as this information is often not reported in many studies. These differences in averaging methods also complicate direct comparisons between works. Please note that there may be additional subtle variations in the experimental setup that are not captured in these tables, which could prevent a fair comparison.
| Micro | Macro | Part. Micro | |||||||||
| Reported In | P | R | F1 | P | R | F1 | P | F1 | Dataset | Method | Experimental Setup |
| Closed Domain Problem Framing | |||||||||||
| Hou2019Leaderboards | 60.2 | 73.1 | 66.0 | 54.1 | 65.9 | 56.6 | NLP-TDMS-v1 | TDMS-IE | |||
| Hou2019Leaderboards | 29.4 | 42.0 | 34.6 | 24.9 | 43.6 | 28.1 | NLP-TDMS-v1 | EL† | |||
| Hou2019Leaderboards | 56.8 | 23.8 | 33.6 | 56.8 | 30.9 | 37.3 | NLP-TDMS-v1 | MLC† | |||
| Hou2019Leaderboards | 16.8 | 7.8 | 10.6 | 8.1 | 6.4 | 6.9 | NLP-TDMS-v1 | SM† | |||
| Hou2019Leaderboards | 60.8 | 76.8 | 67.8 | 62.5 | 75.2 | 65.3 | NLP-TDMS-v2 | TDMS-IE | |||
| Hou2019Leaderboards | 24.3 | 36.3 | 29.1 | 18.1 | 31.8 | 20.5 | NLP-TDMS-v2 | EL† | |||
| Hou2019Leaderboards | 42.0 | 20.9 | 27.9 | 42.0 | 23.1 | 27.8 | NLP-TDMS-v2 | MLC† | |||
| Hou2019Leaderboards | 36.0 | 19.6 | 25.4 | 31.8 | 30.6 | 31.0 | NLP-TDMS-v2 | SM† | |||
| Hou2019Leaderboards | 68.6 | 40.3 | 50.8 | 29.6 | 29.1 | 28.1 | NLP-TDMS-v2 | TDMS-IE | TAE# | ||
| Hou2019Leaderboards | 50.0 | 23.7 | 32.2 | 20.8 | 20.1 | 19.4 | NLP-TDMS-v2 | TDMS-IE | TAT# | ||
| Hou2019Leaderboards | 47.9 | 14.2 | 21.9 | 11.3 | 11.3 | 10.7 | NLP-TDMS-v2 | TDMS-IE | TA# | ||
| Kardas2020AXCELL | 65.8 | 58.5 | 61.9 | 56.0 | 55.8 | 54.1 | NLP-TDMS-v3 | AxCell | |||
| Kardas2020AXCELL | 53.4 | 66.3 | 59.2 | 57.1 | 66.1 | 58.5 | NLP-TDMS-v3 | TDMS-IE | |||
| Kardas2020AXCELL | 67.8 | 47.8 | 56.1 | 47.9 | 46.4 | 43.5 | PwC-LB-v1 | AxCell | |||
| Kabongo2021Mining | 76.4 | 66.4 | 71.1 | 63.5 | 64.1 | 61.4 | NLP-TDMS-v1 | ORKG-TDM | XLNet | ||
| Kabongo2021Mining | 65.3 | 73.1 | 69.0 | 57.6 | 68.7 | 60.1 | NLP-TDMS-v1 | ORKG-TDM | SciBERT | ||
| Kabongo2021Mining | 79.5 | 57.6 | 66.8 | 59.0 | 55.4 | 54.7 | NLP-TDMS-v1 | ORKG-TDM | BERT | ||
| Kabongo2021Mining | 77.1 | 70.9 | 73.9 | 71.7 | 73.9 | 70.6 | NLP-TDMS-v2 | ORKG-TDM | XLNet | ||
| Kabongo2021Mining | 79.6 | 63.3 | 70.5 | 68.1 | 67.5 | 65.5 | NLP-TDMS-v2 | ORKG-TDM | BERT | ||
| Kabongo2021Mining | 65.7 | 76.8 | 70.8 | 65.7 | 77.2 | 68.3 | NLP-TDMS-v2 | ORKG-TDM | SciBERT | ||
| Kabongo2021Mining | 95.1 | 92 | 93.5 | 92.3 | 93.5 | 91.7 | ORKG-PwC-v1 | ORKG-TDM | XLNet TAET# | ||
| Kabongo2021Mining | 93.5 | 93.2 | 93.3 | 90.5 | 94.4 | 91.2 | ORKG-PwC-v1 | ORKG-TDM | XLNet TAT# | ||
| Kabongo2021Mining | 95.0 | 90.5 | 92.7 | 91.6 | 93.1 | 91.2 | ORKG-PwC-v1 | ORKG-TDM | XLNet# | ||
| Kabongo2021Mining | 95.7 | 88.3 | 91.8 | 91.7 | 92.1 | 90.8 | ORKG-PwC-v1 | ORKG-TDM | BERT | ||
| Kabongo2021Mining | 94.2 | 89 | 91.5 | 89.2 | 91.5 | 89.2 | ORKG-PwC-v1 | ORKG-TDM | XLNet TAE# | ||
| Kabongo2021Mining | 94.4 | 87.6 | 90.9 | 89.7 | 91.4 | 89.4 | ORKG-PwC-v1 | ORKG-TDM | SciBERT | ||
| Kabongo2021Mining | 92.6 | 90 | 91.3 | 88.6 | 92.9 | 89.4 | ORKG-PwC-v1 | ORKG-TDM | XLNet TA# | ||
| Kabongo2021Mining | 94.9 | 91.2 | 93.0 | 92.8 | 94.8 | 92.8 | ORKG-PwC-v2 | ORKG-TDM | XLNet | ||
| Kabongo2021Mining | 95.5 | 89.1 | 92.1 | 92.8 | 93.9 | 92.4 | ORKG-PwC-v2 | ORKG-TDM | BERT | ||
| Kabongo2021Mining | 94.1 | 88.5 | 91.2 | 90.9 | 93.4 | 91.1 | ORKG-PwC-v2 | ORKG-TDM | SciBERT | ||
| Kabongo2023ORKG | 95.2 | 92.2 | 93.6 | 91.5 | 93.3 | 91.3 | ORKG-PwC-v5 | ORKG-LB | BigBERT | ||
| Kabongo2023ORKG | 94.8 | 93.9 | 94.3 | 91.3 | 94.4 | 91.8 | ORKG-PwC-v5 | ORKG-LB | BERT | ||
| Kabongo2023ORKG | 94.8 | 93.9 | 94.3 | 91.3 | 94.4 | 91.8 | ORKG-PwC-v5 | ORKG-LB | SciBERT | ||
| Kabongo2023ORKG | 95.4 | 93.9 | 94.7 | 93.2 | 95.7 | 93.5 | ORKG-PwC-v6 | ORKG-LB | BERT | ||
| Kabongo2023ORKG | 95.4 | 91.1 | 93.2 | 92.6 | 94.3 | 92.2 | ORKG-PwC-v6 | ORKG-LB | SciBERT | ||
| Kabongo2023ORKG | 93.2 | 94.9 | 93.0 | 95.7 | 92.4 | 94.0 | ORKG-PwC-v6 | ORKG-LB | BigBERT | ||
| Kabongo2023ORKG | 95.1 | 94.6 | 94.8 | 93.1 | 96.4 | 93.7 | ORKG-PwC-v6 | ORKG-LB | XLNet | ||
| Kabongo2023ORKG | 95.4 | 88.0 | 91.5 | 91.2 | 92.3 | 90.6 | ORKG-PwC-v3 | ORKG-LB | BERT | ||
| Kabongo2023ORKG | 93.7 | 86.0 | 89.7 | 89.4 | 91.7 | 89.2 | ORKG-PwC-v3 | ORKG-LB | SciBERT | ||
| Kabongo2023ORKG | 93.6 | 85.3 | 89.3 | 87.5 | 88.7 | 86.6 | ORKG-PwC-v3 | ORKG-LB | BigBird | ||
| Kabongo2023ORKG | 94.9 | 91.2 | 93.0 | 91.9 | 94.4 | 92.0 | ORKG-PwC-v4 | ORKG-LB | XLNet | ||
| Kabongo2023ORKG | 96.0 | 90.0 | 92.9 | 93.5 | 94.2 | 92.8 | ORKG-PwC-v4 | ORKG-LB | BERT | ||
| Kabongo2023ORKG | 94.6 | 88.6 | 91.5 | 91.7 | 93.9 | 91.6 | ORKG-PwC-v4 | ORKG-LB | SciBERT | ||
| Kabongo2023ORKG | 94.6 | 87.2 | 90.7 | 90.7 | 91.6 | 89.7 | ORKG-PwC-v4 | ORKG-LB | BigBird | ||
| Kabongo2023ZeroShot | 9.2 | 78.1 | 16.5 | 14.3 | 86.6 | 21.9 | ORKG-PwC-v7* | ORKG-TDM | XLNet | ||
| Kabongo2023ZeroShot | 14.1 | 72.9 | 23.6 | 20.1 | 83.4 | 28.9 | ORKG-PwC-v7* | ORKG-TDM | BERT | ||
| Kabongo2023ZeroShot | 10.4 | 81.7 | 18.4 | 16.2 | 89 | 24.4 | ORKG-PwC-v7* | ORKG-TDM | BERT | ||
| Kabongo2023ZeroShot | 10.1 | 76.8 | 17.8 | 14.9 | 86.4 | 22.7 | ORKG-PwC-v7* | ORKG-TDM | XLNet | ||
| sahinuc2024efficient | 55.1 | 25.8 | 35.1 | SciLead | AxCell | ||||||
| sahinuc2024efficient | 40.7 | 39.5 | 40.1 | SciLead | TDMR-PR | Llama 2+CS | |||||
| sahinuc2024efficient | 35.9 | 34.9 | 35.4 | SciLead | TDMR-PR | Llama 2 | |||||
| sahinuc2024efficient | 58.4 | 52.1 | 55.1 | SciLead | TDMR-PR | Mixtral+CS | |||||
| sahinuc2024efficient | 55.7 | 48.8 | 51.0 | SciLead | TDMR-PR | Mixtral | |||||
| sahinuc2024efficient | 62.0 | 58.1 | 60.0 | SciLead | TDMR-PR | Llama 3+CS | |||||
| sahinuc2024efficient | 77.1 | 72.6 | 74.8 | SciLead | TDMR-PR | Llama 3 | |||||
| sahinuc2024efficient | 69.0 | 63.8 | 66.3 | SciLead | TDMR-PR | GPT-4+CS | |||||
| sahinuc2024efficient | 75.3 | 70.4 | 72.8 | SciLead | TDMR-PR | GPT-4 | |||||
| Open Domain Problem Framing | |||||||||||
| Yang2022TELIN | 68.2 | 45.3 | 56.5 | 49.7 | 43.1 | 42.5 | PwC-LB-v2 | TELIN | |||
| Hybrid Domain Problem Framing | |||||||||||
| sahinuc2024efficient | 27.23 | 22.99 | 24.93 | SciLead | TDMR-PR | Llama 2 | |||||
| sahinuc2024efficient | 27.89 | 24.48 | 26.07 | SciLead | TDMR-PR | Mixtral | |||||
| sahinuc2024efficient | 50.75 | 45.30 | 47.87 | SciLead | TDMR-PR | Llama 3 | |||||
| sahinuc2024efficient | 56.08 | 51.89 | 53.90 | SciLead | TDMR-PR | GPT-4 | |||||
| †SM, MLC, and EL are baseline methods, representing String Match, Multi-Label Classification, and Entity Linking, respectively. * trained on ORKG-PwC-v6/v7. # REC, TAET, and Full refer to DocREC, DocTAET, and the Full Paper representations of the document, respectively. These are reported as part of an ablation study examining different document representations. For more details on these representations, see § 5.1. | |||||||||||
| Micro | Macro | Part. Micro | |||||||||
| Reported In | P | R | F1 | P | R | F1 | P | F1 | Dataset | Method | Experimental Setup |
| Closed Domain Problem Framing | |||||||||||
| Hou2019Leaderboards | 10.8 | 13.1 | 11.8 | 9.3 | 11.8 | 9.9 | NLP-TDMS-v1 | TDMS-IE | |||
| Hou2019Leaderboards | 3.8 | 1.8 | 2.4 | 1.3 | 1.0 | 1.1 | NLP-TDMS-v1 | SM† | |||
| Hou2019Leaderboards | 6.8 | 2.9 | 4.0 | 6.8 | 6.1 | 6.2 | NLP-TDMS-v1 | MLC† | |||
| Kardas2020AXCELL | 27.4 | 24.4 | 25.8 | 20.2 | 20.6 | 19.7 | NLP-TDMS-v3 | AxCell | |||
| Kardas2020AXCELL | 6.8 | 8.4 | 7.5 | 8.6 | 9.5 | 8.8 | NLP-TDMS-v3 | TDMS-IE | |||
| Kardas2020AXCELL | 37.4 | 23.2 | 28.7 | 24.0 | 21.8 | 21.1 | PwC-LB-v1 | AxCell | |||
| sahinuc2024efficient | 32.59 | 13.67 | 19.26 | SciLead | AxCell | ||||||
| sahinuc2024efficient | 10.06 | 21.59 | 13.73 | SciLead | TDMR-PR | Llama 2+CS | |||||
| sahinuc2024efficient | 9.63 | 15.25 | 11.81 | SciLead | TDMR-PR | Llama 2 | |||||
| sahinuc2024efficient | 26.54 | 24.61 | 25.54 | SciLead | TDMR-PR | Mixtral+CS | |||||
| sahinuc2024efficient | 24.66 | 21.73 | 23.10 | SciLead | TDMR-PR | Mixtral | |||||
| sahinuc2024efficient | 23.22 | 29.54 | 26.00 | SciLead | TDMR-PR | Llama 3+CS | |||||
| sahinuc2024efficient | 27.11 | 35.60 | 30.78 | SciLead | TDMR-PR | Llama 3 | |||||
| sahinuc2024efficient | 49.82 | 48.71 | 49.26 | SciLead | TDMR-PR | GPT-4+CS | |||||
| sahinuc2024efficient | 56.02 | 54.53 | 55.27 | SciLead | TDMR-PR | GPT-4 | |||||
| Open Domain Problem Framing | |||||||||||
| Yang2022TELIN | 38.3 | 20.8 | 26.3 | 26.6 | 19.2 | 21.3 | PwC-LB-v2 | TELIN | |||
| Hybrid Domain Problem Framing | |||||||||||
| sahinuc2024efficient | 4.17 | 9.89 | 5.87 | SciLead | TDMR-PR | Llama 2 | |||||
| sahinuc2024efficient | 14.65 | 12.27 | 13.35 | SciLead | TDMR-PR | Mixtral | |||||
| sahinuc2024efficient | 15.70 | 18.75 | 17.09 | SciLead | TDMR-PR | Llama 3 | |||||
| sahinuc2024efficient | 40.60 | 39.56 | 40.07 | SciLead | TDMR-PR | GPT-4 | |||||
| sahinuc2024efficient | 51.01 | 51.03 | 51.02 | SciLead | TDMR-PR | GPT-4 FS | |||||
| †SM, MLC, and EL are baseline methods, representing String Match, Multi-Label Classification, and Entity Linking, respectively. | |||||||||||
| Micro | Macro | Part. Micro | |||||||||
| Reported In | P | R | F1 | P | R | F1 | P | F1 | Dataset | Method | Experimental Setup |
| Closed Domain Problem Framing | |||||||||||
| Jain2020SCIREX | 0.48 | 0.89 | 0.62 | SciREX | TDMS-IE | ||||||
| Open Domain Problem Framing | |||||||||||
| Jain2020SCIREX | 0.53 | 0.72 | 0.61 | SciREX | SciREX-IE | ||||||
| Micro | Macro | Part. Micro | |||||||||
| Reported In | P | R | F1 | P | R | F1 | P | F1 | Dataset | Method | Experimental Setup |
| Closed Domain Problem Framing | |||||||||||
| Kardas2020AXCELL | 70.6 | 57.3 | 63.3 | 60.7 | 62.6 | 59.7 | PwC-LB-v1 | AxCell | |||
| Kabongo2021Mining | 97.4 | 93.6 | 95.5 | 93.7 | 94.8 | 93.6 | ORKG-PwC-v1 | ORKG-TDM | XLNet | ||
| Kabongo2023ORKG | 96.8 | 95.9 | 96.4 | 94.3 | 97.2 | 95.0 | ORKG-PwC-v6 | ORKG-LB | XLNet | ||
| Kabongo2023ORKG | 96.8 | 95.9 | 96.4 | 94.3 | 97.2 | 95.0 | ORKG-PwC-v4 | ORKG-LB | XLNet | ||
| sahinuc2024efficient | 68.98 | 58.52 | 63.32 | SciLead | AxCell | ||||||
| sahinuc2024efficient | 59.83 | 67.20 | 63.30 | SciLead | TDMR-PR | Llama 2+CS | |||||
| sahinuc2024efficient | 55.45 | 60.74 | 57.97 | SciLead | TDMR-PR | Llama 2 | |||||
| sahinuc2024efficient | 86.27 | 91.99 | 89.04 | SciLead | TDMR-PR | Mixtral+CS | |||||
| sahinuc2024efficient | 86.85 | 89.74 | 88.27 | SciLead | TDMR-PR | Mixtral | |||||
| sahinuc2024efficient | 85.69 | 90.85 | 88.19 | SciLead | TDMR-PR | Llama 3+CS | |||||
| sahinuc2024efficient | 87.33 | 92.17 | 89.68 | SciLead | TDMR-PR | Llama 3 | |||||
| sahinuc2024efficient | 90.70 | 90.77 | 90.73 | SciLead | TDMR-PR | GPT-4+CS | |||||
| sahinuc2024efficient | 90.62 | 91.10 | 90.86 | SciLead | TDMR-PR | GPT-4 | |||||
| Open Domain Problem Framing | |||||||||||
| Yang2022TELIN | 70.3 | 53.7 | 59.2 | 60.5 | 57.3 | 57.1 | PwC-LB-v2 | TELIN | |||
[…]